# Dynamics of binding three independent ligands to a single scaffold

### v3

@@ 3 Detailed balance
and similarly for all other complexes. Thus, at steady state, the conservation laws can be rewritten, after changing to new variables $a=K_{\mbox{\tiny A}}a$, $b=K_{\mbox{\tiny B}}b$, $c=K_{\mbox{\tiny C}}c$, as follows:
and we will call these the polynomial form of the conservation laws. They represent a set of four polynomial equations in four variables, and the $(y,a,b,c)$ coordinates of the steady state constitute their unique positive solution. Conversely, the solution of these equations gives rise to the steady state of the original system once that we define $yab:=y\cdot a\cdot b$ and so forth (recall that we are assuming variables have been rescaled so that all $K_{\mbox{\tiny L}}=1$).
@@ 5 Steady state of YABC as a function of $Y_{\text{tot}}$
$f(r)\;<\;A_{\text{tot}}B_{\text{tot}}C_{\text{tot}}\frac{y(r)}{(1+y(r))^{3}}\;<\;A_{\text{tot}}B_{\text{tot}}C_{\text{tot}}\frac{1+y(r)}{(1+y(r))^{3}}\;=\;A_{\text{tot}}B_{\text{tot}}C_{\text{tot}}\frac{1}{(1+y(r))^{3}}$ (5)
@@
Figure 3: Example of a typical function $f(r)$. The parameters here are: $(A_{\text{tot}},B_{\text{tot}},C_{\text{tot}})=(1,5,10)$ and (without loss of generality, after rescaling) $K_{\mbox{\tiny A}}=K_{\mbox{\tiny B}}=K_{\mbox{\tiny C}}=1$.
@@ 5.2 The bispecific case
Existence and uniqueness follow by the same argument as in the trispecific case (or could be proved directly). Smoothness follows by the implicit function theorem, as the Jacobian of the left-hand side has determinant $(1+a_{1}+y+ya_{2})(1+a_{2}+y+ya_{1})$, which is nonzero (in fact, positive) for all $(y,a_{1},a_{2})$. As in the trispecific case, $\lim_{r\rightarrow 0}f(r)=\lim_{r\rightarrow\infty}f(r)=0$. Thus we only need to show that the derivative $f^{\prime}(r)$ can vanish at most once.

### v4

@@ 3 Detailed balance
and similarly for all other complexes.
@@ 5 Steady state of YABC as a function of $Y_{\text{tot}}$
We will assume for simplicity that $K_{\mbox{\tiny A}}=K_{\mbox{\tiny B}}=K_{\mbox{\tiny C}}=1$ and leave the general case for a revised version. Thus, at steady state, the conservation laws can be rewritten as follows:
and we will call these the polynomial form of the conservation laws. They represent a set of four polynomial equations in four variables, and the $(y,a,b,c)$ coordinates of the steady state constitute their unique positive solution. Conversely, the solution of these equations gives rise to the steady state of the original system once that we define $yab:=y\cdot a\cdot b$ and so forth.
@@
$f(r)\;<\;A_{\text{tot}}B_{\text{tot}}C_{\text{tot}}\frac{y(r)}{(1+y(r))^{3}}\;<\;A_{\text{tot}}B_{\text{tot}}C_{\text{tot}}\frac{1+y(r)}{(1+y(r))^{3}}\;=\;A_{\text{tot}}B_{\text{tot}}C_{\text{tot}}\frac{1}{(1+y(r))^{2}}$ (5)
@@
Figure 3: Example of a typical function $f(r)$. The parameters here are: $(A_{\text{tot}},B_{\text{tot}},C_{\text{tot}})=(1,5,10)$ and $K_{\mbox{\tiny A}}=K_{\mbox{\tiny B}}=K_{\mbox{\tiny C}}=1$.
@@ 5.2 The bispecific case
Existence and uniqueness follow by the same argument as in the trispecific case (or could be proved directly). (See also [11, 4] for other proofs, including non-detailed balanced cases.) Smoothness follows by the implicit function theorem, as the Jacobian of the left-hand side has determinant $(1+a_{1}+y+ya_{2})(1+a_{2}+y+ya_{1})$, which is nonzero (in fact, positive) for all $(y,a_{1},a_{2})$. As in the trispecific case, $\lim_{r\rightarrow 0}f(r)=\lim_{r\rightarrow\infty}f(r)=0$. Thus we only need to show that the derivative $f^{\prime}(r)$ can vanish at most once.
@@
Remark 5.4.
One can show that the following functions of $(r,p_{1},p_{2})$ solve the equations (8)-(10). For convenience, introduce
$P_{1}=\sqrt{(r-p_{1}-1)^{2}+4r}$
and
$P_{2}=\sqrt{(r-p_{2}-1)^{2}+4r}\,.$
Then, at steady state:
$y$ $=$ $\frac{(p_{1}-r+1-P_{1})\,(p_{2}-r+1-P_{2})}{4r}$
$a_{1}$ $=$ $\frac{p_{1}-r-1+P_{1}}{2}$
$a_{2}$ $=$ $\frac{p_{2}-r-1+P_{2}}{2}\,.$
These formulas can be verified by substitution. Furthermore, one can algebraically show, after simplification, that
$y\cdot a_{1}\cdot a_{2}\;=\;\frac{1}{4r}\,(r+p_{1}+1-P_{1})(r+p_{2}+1-P_{2}).$
This last formula can be easily shown to be the same as the one obtained in [4], when $K_{\mbox{\tiny A}}=K_{\mbox{\tiny B}}=1$. (The formula for $y$ would match, after simplifications, eqn. (S57) (for “[B]”) in that reference but only if there is an $r$ instead of $r^{2}$ in the denominator.)